# Consequences of a Two-Time Relativistic Bohmian Model

### v16

@@ IV.1 Zitterbewegung in arbitrary direction
With respect to an arbitrary spatial direction $\hat{s}\neq\hat{v}_{c}$, let’s consider the polar coordinate system ($s$, $\theta$, $\phi$), where $\theta\neq 0$ is the angle between $\hat{s}$ and $\hat{v}_{c}$ and $\phi$ the azimuth (see next figure).
@@
$\vec{\nabla}R=\frac{dR}{ds}\hat{s}\quad\quad\nabla^{2}R=\frac{2}{s}\frac{dR}{ds}+\frac{d^{2}R}{ds^{2}}$ (49)
@@
$\frac{\partial^{2}R}{\partial t^{2}}-{v_{c_{s}}^{2}}R^{\prime\prime}=0$ (50)
where $v_{c_{s}}=v_{c}\ cos\theta$ and the apostrophe indicates derivative over $s$. Substituting eq. (50) and the second of eqs. (49) into eq. (11) and rooting is found:
$H=\sqrt{m^{2}\gamma_{c}^{2}c^{4}-\hbar^{2}\frac{c^{2}}{\gamma_{c_{s}}^{2}}\frac{R^{\prime\prime}}{R}+\frac{2\hbar^{2}}{sR}R^{\prime}}\ ,\ \gamma_{c_{s}}\neq\gamma_{c}$ (51)
@@
$v_{c_{s}}\,\frac{R^{\prime}}{R}\,(m\,\gamma_{c}\,c^{2}-H)+\frac{1}{2}\frac{\partial H}{\partial t}=0$ (52)
@@
The different value of $\lambda$ therefore causes a dependence of the oscillation frequency on the spatial direction: this is precisely the relativistic anisotropy $v_{i_{s}}\neq v_{i_{u}}$ that we anticipated in section III. Remarkably, this is reflected in a spatial anisotropy of the uncertainty principle, where we have the presence of $\gamma_{c_{s}}$ instead of $\gamma_{c}$ to divide:
$\Delta s\cdot\Delta p_{s}\sim\frac{\hbar}{\gamma_{c_{s}}}$ (53)
@@ IV.3 Standard Zitterbewegung
Interpretating the generic oscillatory motion in $\tau$ as occurring in t, the standard Zitterbewegung is re-obtained. In the standard view, $R$ is considered time independent, so the second side of eq. (10) is just $-\frac{H^{2}}{\hbar{}^{2}}$, with $H=const$. In the new theory this means demanding: $\frac{1}{R}\frac{\partial^{2}R}{\partial t^{2}}-2\,\frac{i}{\hbar}\,\frac{H}{R}\frac{\partial R}{\partial t}=0$, that provides $R\propto\,e^{\frac{2i}{\hbar}H\,t}$ and hence $V_{Q}\propto e^{-\frac{4i}{\hbar}H\,t}$. Then, using eqs. (22) for $V_{Q}$ and (24) (both as for t-time), after having conveniently determined the arbitrary constants, one can get: $v_{i_{s}}\propto c\,e^{-\frac{2i}{\hbar}H\,t}$, as found in the Heisenberg picture.
@@ V.1 Collapses
The apparent observational ”collapse” is an extreme case of destructive interference that occurs after a energy variation of the particle for collision (more generally, interaction) with an external particle or field: what happens is that the self-wavefunction at the point of impact changes abruptly its phase proportional to $S_{t}$; then, its interference with the self-wavefunctions located elsewhere cancels - instantaneously, with respect to $t$-time - the latter. From a corpuscular point of view, the Quantum Potential no longer allows the particle to $\tau$-move away from the proximity of the impact.
@@
For what has been said, the paradox that the particle, being $t$-simultaneously in every point of a certain macroscopic volume, violates the conservation of mass and of number of particles is only apparent: a particle search experiment will always detect only one particle. The total purely quantum energy distributed in the sphere-like volume at any $t$-instant is indeed infinite, but it is beyond any kind of experimentation. Actually it is virtual, in the same sense used in standard view. Indeed, the existence of the hidden time $\tau$ is capable of explaining why the increasingly larger energies predicted by entering the infinitesimal are real (as experiments excellently confirm) but never directly observed.
@@ VI Discussion
In the explored theory, the apparently paradoxical aspects of Quantum Mechanics are just the effect of the existence of an extra unobservable time dimension. By traveling in a hidden time, you can deterministically explain why a particle or photon is able to interact instantly with more slits in a diffraction, to run across multiple paths between two points, to admit entanglements with other arbitrarily distant particles, to emit and reabsorb many species of virtual particles… All established facts.
@@
$\frac{\partial R^{2}}{\partial\tau}-v_{i_{s}}{\frac{dR^{2}}{ds}}=0$ (54)
When one refers this motion to an oscillation, one realizes that $R^{2}$ is in antiphase - i.e. shifted by half a period - with respect to the particle moving with speed $v_{i_{s}}$. This ”pursuing” reminds the way the Zitterbewegung is interpreted in the standard interpretation: the particle ”chases” its antiparticle (and vice versa) which, according to the uncertainty principle, is created for lengths $\sim\lambda_{c}$ (see e.g. [71]). However, our model makes no reference to antimatter, admitting only positive values for energy. In fact, we are able to justify this equation by referring exclusively to the meaning of the physical quantities involved. According to our model, the operation of finding the particle in a unit volume (whose probability is $R^{2}$) must specify a $\bar{t}$-event. Then, for an amplitude $\sim\lambda_{c}$, $\bar{t}$ has an uncertainty of about half a period of the intrinsic oscillation of the particle, as we have shown. And this effectively means that the particle must perform half an oscillation before the measurement is achievable, ending up in antiphase with $R^{2}$.

### v18

@@ IV.1 Zitterbewegung in arbitrary direction
With respect to an arbitrary spatial direction $\hat{s}\neq\hat{v}_{c}$, let’s consider the polar coordinate system ($s$, $\theta$, $\phi$), where $\theta\neq 0$ is the angle between $\hat{s}$ and $\hat{v}_{c}$ and $\phi$ the azimuth (fig. 1).
Figure 1: Polar coordinate system for studying the Zitterbewegung in an arbitrary direction $s$.
@@
$\vec{\nabla}R=\frac{{\rm{d}}R}{{\rm{d}}s}\hat{s},\quad\quad\nabla^{2}R=\frac{2}{s}\frac{{\rm{d}}R}{{\rm{d}}s}+\frac{{\rm{d}}^{2}R}{{\rm{d}}s^{2}}.$ (49)
@@
$\frac{\partial^{2}R}{\partial t^{2}}-{v_{c_{s}}^{2}}R^{\prime\prime}=0,$ (50)
where $v_{c_{s}}=v_{c}\ {\rm{cos}}\ \theta$ and the apostrophe indicates derivative over $s$. Substituting eq. (50) and the second of eqs. (49) into eq. (11) and rooting is found:
$H=\sqrt{m^{2}\gamma_{c}^{2}c^{4}-\hbar^{2}\frac{c^{2}}{\gamma_{c_{s}}^{2}}\frac{R^{\prime\prime}}{R}-\frac{2\hbar^{2}c^{2}}{sR}R^{\prime}}\ ,\ \gamma_{c_{s}}\neq\gamma_{c},$ (51)
@@
$v_{c_{s}}\,\frac{R^{\prime}}{R}\,(m\,\gamma_{c}\,c^{2}-H)+\frac{1}{2}\frac{\partial H}{\partial t}=0.$ (52)
@@
The different value of $\lambda$ therefore causes a dependence of the oscillation amplitude and frequency on the spatial direction: this is precisely the relativistic anisotropy $v_{i_{s}}\neq v_{i_{u}}$ that we anticipated in section III. Remarkably, this is reflected in a spatial anisotropy of the uncertainty principle, where we have the presence of $\gamma_{c_{s}}$ instead of $\gamma_{c}$ to divide:
$\Delta s\cdot\Delta p_{s}\sim\frac{\hbar}{\gamma_{c_{s}}}.$ (53)
@@ IV.3 Standard Zitterbewegung
Interpretating the generic oscillatory motion in $\tau$ as occurring in t, the standard Zitterbewegung is re-obtained. In the standard view, $R$ is considered time independent, so the second side of eq. (10) is just $-\frac{H^{2}}{\hbar{}^{2}}$, with $H=const$. In the new theory this means demanding: $\frac{1}{R}\frac{\partial^{2}R}{\partial t^{2}}-2\,\frac{\rm{i}}{\hbar}\,\frac{H}{R}\frac{\partial R}{\partial t}=0$, that provides $R\propto\,{\rm{e}}^{\frac{2{\rm{i}}}{\hbar}H\,t}$ and hence $V_{Q}\propto{\rm{e}}^{-\frac{4{\rm{i}}}{\hbar}H\,t}$. Then, using eqs. (22) for $V_{Q}$ and (24) (both as for t-time), after having conveniently determined the arbitrary constants, one can get: $v_{i_{s}}\propto c\,{\rm{e}}^{-\frac{2{\rm{i}}}{\hbar}H\,t}$, as found in the Heisenberg picture.
@@ V.1 Collapses
The apparent observational ”collapse” is an extreme case of destructive interference that occurs after a energy variation of the particle for collision (more generally, interaction) with an external particle or field: what happens is that the packet wavefunctions located in the impact area abruptly change their phase proportional to $S_{t}$ and then interfere destructively (statistically) with those located elsewhere. Hence the packet reconfigures itself - instantaneously, with respect to $t$-time - by centering itself in the impact area and restricting its dispersion. From a corpuscular point of view, the Quantum Potential no longer allows the particle to $\tau$-move away from the proximity of the impact.
@@
For what has been said, the paradox that the particle, being $t$-simultaneously in every point of a certain macroscopic volume, violates the conservation of mass and of number of particles is only apparent: a particle search experiment will always detect only one particle. The total purely quantum energy distributed in the sphere-like volume at any $t$-instant is indeed infinite, but it is beyond any kind of direct experimentation. Actually it is virtual, in the same sense used in standard view. Indeed, the existence of the hidden time $\tau$ is capable of explaining why the increasingly larger energies predicted by entering the infinitesimal are real (as experiments excellently confirm) but never directly observed.
@@ VI Discussion
In the explored theory, the non-locality and the apparently paradoxical aspects of Quantum Mechanics are just the effect of the existence of an extra unobservable time dimension. By traveling in a hidden time, you can deterministically explain why a particle or photon is able to interact instantly with more slits in a diffraction, to run across multiple paths between two points, to admit entanglements with other arbitrarily distant particles, to emit and reabsorb many species of virtual particles… All established facts.
@@
$\frac{\partial R^{2}}{\partial\tau}-v_{i_{s}}{\frac{{\rm{d}}R^{2}}{{\rm{d}}s}}=0.$ (54)
When one refers this motion to an oscillation, one realizes that $R^{2}$ is in antiphase - i.e. shifted by half a period - with respect to the particle moving with speed $v_{i_{s}}$. This ”pursuing” reminds the way the Zitterbewegung is interpreted in the standard interpretation: the particle ”chases” its antiparticle (and vice versa) which, according to the uncertainty principle, is created for lengths $\sim\lambda_{c}$ (see e.g. [71]). However, our model makes no reference to antimatter, admitting only positive values for energy. In fact, we are able to justify this equation by referring exclusively to the meaning of the physical quantities involved. According to our model, the operation of finding the particle in a unit volume (whose probability is $R^{2}$) must specify a $\bar{t}$-event. Then, for an amplitude $\sim\lambda_{c}$, $\bar{t}$ must have an uncertainty of about half a period of the intrinsic oscillation of the particle, as we have shown. And this effectively means that the particle must perform half an oscillation - the cause of uncertainty - before the measurement is achievable, ending up in antiphase with $R^{2}$.
VII Acknowledgments
I am indebted to H. Nikolić and A. J. Silenko for their support. I also especially thank the reviewer of the manuscript for encouraging me to improve it significantly. Besides, the promoters and developers of WolframAlpha to make free available on the Web their computational engine.